# Herd Behaviour in Public Goods Games

### v2

@@ Abstract
The problem of free-riding arises when individuals benefit from a shared resource, service, or public good without proportionately contributing to its provision. This conduct often leads to a collective action problem, as individuals pursue personal gains while relying on the contributions of others. In this study, we present a Bayesian inference model to elucidate the behaviour of participants in a Public Goods Game, a conceptual framework that captures the essence of the free-riding problem. Here, individuals possess information on the distribution of group donations to the public good. Our model is grounded in the premise that individuals strive to harmonise their actions with the group’s donation patterns. Our model is able to replicate behavioural patterns that resemble those observed in experiments with mid-sized groups (100 people) but fails to replicate those for larger scales (1000 people). These insights contribute to the understanding of how cooperation is achieved through alignment with group behaviour.
@@ I Introduction
When exploring the complex dynamics of human cooperation and decision-making, the provision of public goods has attracted a great deal of attention in recent decades. At the core of this research lies the profound dilemma between individually rational decisions and the collective optimal outcome. The extensively researched paradigm for this issue is the Public Goods Game (PGG) [1], where $N\geq 2$ agents can voluntarily contribute to a common pool. The pool is then multiplied or enhanced by a factor $1<r<N$, and the resulting value is shared among all participants, regardless of their individual contributions. The Nash Equilibrium in this game is characterised by investing zero amount in the public good, while the collective optimal outcome demands maximal contributions from all participants. Beyond its theoretical intrigue, this dilemma is of vital importance in addressing real-world challenges such as climate change mitigation, public infrastructure management, ecosystem protection, and sustainable exploitation, where solutions hinge on voluntary contributions from many. Despite game theory predictions ([2, 3, 4], experimental findings, both in controlled laboratory settings and online, consistently show that participants contribute approximately half of their endowment in one-shot games; and in iterated games, contributions decline but persist, with a general pattern of gradual decrease as players learn the consequences of their actions. The observed results are influenced by the information that participants receive about the actions of other players [5, 6, 7, 8, 9, 10].
@@
Much research has explored individual contributions in traditional Public Goods Games, where decisions hinge on the average contributions of the group members. Studies conducted by [11, 12, 13] convincingly showcase the widespread occurrence of conditional cooperation in the Public Goods game. Using the strategy method, participants were categorised into two predominant groups: conditional cooperators, who are inclined to contribute more as others contribute, and free-riders, who abstain from contributions regardless of others’ behaviour. In conditions where individuals cannot base their decisions on others’ behaviour, conditional cooperators, as demonstrated by  [11, 12], determine their contributions relying on diverse beliefs about the contributions of other group members. As far as we are aware, only one study [10] has studied human decisions when presented with the distribution of decisions (not solely the average contribution of the group), and to date, no models have explained the behaviour observed in such a setup. Most of the modelling efforts have focused on understanding the group dynamics based on the average contribution to the group. Our model diverges from those standard cooperation models (see [14, 15] for two recent reviews on cooperation), from theoretical modelling approaches [16, 17, 18, 19], and models using the conditional cooperation mechanism [11, 20, 21]. In our proposal, agents base their decisions on the distribution of decisions of the group, employing Bayesian inference to converge toward the group’s patterns of contribution behaviour, as individuals possess pronounced inclinations to conform and mimic the group [22, 23].
@@ II MODEL DYNAMICS
In this section, we present four models for the observed human behaviour under a Public Goods Game in which people are informed of the distribution of contributions of the others. Then, agent donations are based on Bayesian inference. The first model (A) uses a non-informative prior for the agents, the second model (B) uses a data-based prior, the third one (C) includes non-Bayesian agents and uses the data-based prior from (B), and the fourth one (D) includes non-Bayesian agents and the non-informative prior from A.
@@ II.2 Model with Bayesian agents and data-based prior
In this second model, we assume the individuals possess knowledge about the behaviour of people in this kind of dilemmas, and so can estimate a prior distribution of donations to use in their first decision. This would be in line with the observed first donations and donations in one-shot games, which happen to be approximately half of the endowment for a great variety of games (see [29] for a review on the topic). In the model, we use empirical data from [10] which can be found in the Zenodo repository that accompanies the referred paper. The prior distribution is then a Dirichlet distribution with parameters $\alpha_{1}=0.1$, $\alpha_{2}=0.2$, $\alpha_{3}=0.32$, $\alpha_{4}=0.2$, $\alpha_{5}=0.07$, and $\alpha_{6}=0.11$ (Please review the code that accompanies this research.).
@@ II.3 Model with Bayesian agents, data-based prior, and non-Bayesian agents
In experiments with real people, we find a percentage of people who contribute the whole endowment (known as full cooperators) and another percentage that contribute nothing (known as full defectors or free riders).
In this third model, we incorporate agents that implement both types of behaviours. We estimate the percentage of full cooperators and full defectors from the data in  [10]. To do so, we define the full cooperator as someone who donated the whole endowment in more than 80% rounds, and analogously we define the full defector as those people who donated zero in more than 80% of the rounds. Then, we compute the average number of full cooperators and full defectors in the experimental data, resulting in a 5% of full cooperators and 6% of full defectors (for details, see the code accompanying this work).
@@ II.4 Model with Bayesian agents, non-informative prior, and non-Bayesian agents
This last model incorporates the non-informative prior from model A and includes the same percentage of non-Bayesian agents in model C.
@@ III.1 Results for the basic model with Bayesian agents and noninformative prior
Cluster 1 comprises simulations in which the most common cooperative behaviour is centred around the average contribution of the group, along with simulations in which the majority of contributions are above the mean endowment (5 points). See Fig. 3 as a reference. In Cluster 2, simulations are notable for the presence of a high percentage of agents contributing low values of the endowment, with a non-negligible number of simulations featuring a very high percentage of agents who do not contribute at all (Fig. 4). Cluster 4 groups simulations in which the predominant contribution is 20% of the endowment (2 points) (see Supplementary Figure S6). The remaining clusters encapsulate less frequent behaviours (see Supplementary Figures S5, S7 and S8).
@@ III.3 Results for the model with Bayesian agents, data-based prior, and non-Bayesian agents
The results of the cluster analysis with N=100 return 4 clusters (86%, 6%, 2%, and 6%, respectively), shown in Supplementary Figures S21 to S24. The behaviours are grouped into a main cluster that contains 86% of the simulations. These show agent behaviour that becomes more polarised as the rounds of the game progress towards highly polarised values (most agents donate all or nothing at the end of the game), see Fig. 9 as an example. Cluster 2 groups simulations with ultra-cooperative behaviour, i.e. those with an extremely high percentage of agents donating all the endowment (Fig. S22). The remaining clusters, which are less representative, show polarised simulations but to a lesser extent than the previous ones.
@@ III.4 Results for the model with Bayesian agents, non-informative prior, and non-Bayesian agents
The results of the cluster analysis with N=100 indicate the presence of six clusters, representing 82%, 5%, 9%, 1%, 1%, and 2%, respectively. Supplementary Figures S31 to S36 provide a visual representation of these clusters. The majority cluster, with 82% of the simulations, shows ultra-polarised behaviour, due to the presence of free riders and full cooperators (see Fig 11 and S31). With a non-informative prior, the dynamics of the agents are attracted by agents who do not change their opinion. The remaining clusters have simulations that are similar to cluster 1, with some particularities. They can be seen in Supplementary Figures S32 to S36.
Cluster analysis carried out with N=1000 delineates clusters with very similar instances, so we can consider that all behaviours have homogenised. Figure 12 and Supplementary Figure S37 visually illustrate the characteristics of this unique group.
@@ IV Discussion and conclusions
In this study, we have introduced four models to elucidate human behaviour in the Public Goods Game when individuals possess more information than just the group’s average contribution, incorporating details about the distribution of donations from all participants in the previous round. Grounded on the assumption that individuals in this scenario are not seeking to optimise their gains, but rather to converge and imitate the group’s behaviour. The first model, model A, is the simplest, assuming that individuals have no knowledge of how others behave in this game. Model B refines this assumption, incorporating some form of knowledge, drawing insights from real experimental data [10]. The last two models extend these initial models by introducing individuals who are not influenced by the group but exhibit independent behaviours such as free riders or full cooperators, as observed in experiments.
@@
Model C with 100 agents exhibits behavioural patterns compatible with those experimentally observed in groups of 100 individuals (see Figures 1 and 1 vs 9 and Supplementary Figure S21). We observe how agent behaviour and human behaviour in the initial rounds stem from a prior assumption and subsequently evolve, polarising towards situations with a high percentage of full cooperators and free riders.
These findings could be attributed not only to the effect of Bayesian agents but also to a small-scale effect. We note that for larger group sizes, as seen in the results with 1000 agents presented throughout the study, agent behaviours become much more homogeneous across all models.
Our results lead us to hypothesise that humans, in such scenarios, might prefer imitation and convergence behaviours over profit optimisation, providing a plausible explanation for the surprising results, high levels of average cooperation and polarisation, in Public Goods Dilemmas of this kind (individuals with more information).
Adhering to Occam’s Razor, our models, while limited, offer generic insights into how cooperation is achieved through the convergence to group behaviour. A potential extension of the models would be introducing noise or trembling hand in decisions, as found in [28], where ’people are prone to make errors and random decisions’. This avenue presents an intriguing prospect for further exploration.

### v3

@@ Abstract
The problem of free-riding arises when individuals benefit from a shared resource, service, or public good without contributing proportionately to its provision. This conduct often leads to a collective action problem, as individuals pursue personal gains while relying on the contributions of others. In this study, we present a Bayesian inference model to elucidate the behaviour of participants in a Public Goods Game, a conceptual framework that captures the essence of the free-riding problem. Here, individuals possess information on the distribution of group donations to the public good. Our model is grounded in the premise that individuals strive to harmonise their actions with the group’s donation patterns. Our model is able to replicate behavioural patterns that resemble those observed in experiments with midsized groups (100 people), but fails to replicate those for larger scales (1000 people). Our results suggest that, in these scenarios, humans prefer imitation and convergence behaviours over profit optimisation. These insights contribute to understanding how cooperation is achieved through alignment with group behaviour.
@@ I Introduction
When exploring the complex dynamics of human cooperation and decision-making, the provision of public goods has attracted a great deal of attention in recent decades. At the core of this research lies the profound dilemma between individually rational decisions and the collective optimal outcome. The extensively researched paradigm for this issue is the Public Goods Game (PGG) [1], where $N\geq 2$ agents can voluntarily contribute to a common pool. The pool is then multiplied or enhanced by a factor $1<r<N$, and the resulting value is shared among all participants, regardless of their individual contributions. The Nash Equilibrium in this game is characterised by investing zero amount in the public good, while the collective optimal outcome demands maximal contributions from all participants. Beyond its theoretical intrigue, this dilemma is of vital importance in addressing real-world challenges such as climate change mitigation, public infrastructure management, ecosystem protection, and sustainable exploitation, where solutions hinge on voluntary contributions from many. Despite game theory predictions ([2, 3, 4], experimental findings, both in controlled laboratory settings and online, consistently show that participants contribute approximately half of their endowment in one-shot games; and, in iterated games, contributions decline but persist, with a general pattern of gradual decrease as players learn the consequences of their actions. The observed results are influenced by the information that participants receive about the actions of other players [5, 6, 7, 8, 9, 10].
@@
Much research has explored individual contributions in traditional Public Goods Games, where decisions are made after being informed of the last average contribution of the group members. Studies conducted by [11, 12, 13] convincingly showcase the widespread occurrence of conditional cooperation in the Public Goods game. Using the strategy method, participants were categorised into two predominant groups: conditional cooperators, who are inclined to contribute more as others contribute, and free riders, who abstain from contributions regardless of others’ behaviour. In conditions where individuals cannot base their decisions on others’ behaviour, conditional cooperators, as demonstrated by  [11, 12], determine their contributions relying on diverse beliefs about the contributions of other group members. As far as we are aware, only one study [10] has studied human decisions when presented with the distribution of decisions (not solely the average contribution of the group), and to date, no models have explained the behaviour observed in such a setup. Most of the modelling efforts have focused on understanding the group dynamics based on the average contribution to the group. Our model diverges from those standard cooperation models (see [14, 15] for two recent reviews on cooperation), from theoretical modelling approaches [16, 17, 18, 19], and models using the conditional cooperation mechanism [11, 20, 21]. In our proposal, agents base their decisions on the distribution of decisions of the group, employing Bayesian inference to converge toward the group’s patterns of contribution behaviour, as individuals possess pronounced inclinations to conform and mimic the group [22, 23].
@@ II MODEL DYNAMICS
In this section, we present a model for the observed human behaviour under a Public Goods Game in which people are informed of the distribution of contributions of others. In the present model, agent donations are based on Bayesian inference using this information.
@@
The model is implemented in four versions, with different parametrizations. Models (A) and (B) differ on the initial conditions: model (A) uses a non-informative prior for the agents, and model (B) uses a data-based prior. The last two versions are motivated by the present of “special individuals” in real experimental settings, i.e., there is always a percentage of people who do not respond to external feedback or information but always behave as full cooperator or free rider (they donate all or nothing, respectively). We include this more realistic composition of agents to the model. Consequently, these third and fourth models derive from models (A) and (B) including non-Bayesian agents that behave as full cooperators or free riders. Then, the third model (C) includes non-Bayesian agents and uses the data-based prior from (B), and the fourth one (D) includes non-Bayesian agents and the non-informative prior from (A).
@@ II.2 Model with Bayesian agents and data-based prior
In this second model, we assume the individuals possess knowledge about the behaviour of people in this kind of dilemmas, and so can estimate a prior distribution of donations to use in their first decision. This would be in line with the observed first donations and donations in one-shot games, which happen to be approximately half of the endowment for a great variety of games (see [29] for a review on the topic). In the model, we use empirical data from [10] which can be found in the Zenodo repository that accompanies the referred paper. The prior distribution is then a Dirichlet distribution with parameters $\alpha_{1}=0.1$, $\alpha_{2}=0.2$, $\alpha_{3}=0.32$, $\alpha_{4}=0.2$, $\alpha_{5}=0.07$, and $\alpha_{6}=0.11$ (see details in the code that accompanies this manuscript).
@@ II.3 Model with Bayesian agents, data-based prior, and non-Bayesian agents
In this third version, we incorporate non-Bayesian agents: full cooperators and free riders. We estimate the percentage of full cooperators and free riders from the data in  [10]. To do so, we define full cooperator as someone who donated the whole endowment in more than 80% rounds, and, analogously, we define the free rider as those people who donated zero in more than 80% of the rounds. Then, we compute the average number of full cooperators and free riders in the experimental data, resulting in a 5% of full cooperators and 6% of free riders (for details, see the code accompanying this work).
@@ II.4 Model with Bayesian agents, non-informative prior, and non-Bayesian agents
This last model incorporates the non-informative prior from model (A) and includes the same percentage of non-Bayesian agents in model (C).
@@ III.1 Results for the basic model with Bayesian agents and noninformative prior
Cluster 1 comprises simulations in which the most common cooperative behaviour is centred around the average contribution of the group, along with simulations in which the majority of contributions are above the mean endowment (5 points); see Fig. 3 as a reference. In Cluster 2, simulations are notable for the presence of a high percentage of agents contributing low values of the endowment, with a non-negligible number of simulations featuring a very high percentage of agents who do not contribute at all (Fig. 4). Cluster 4 groups simulations in which the predominant contribution is 20% of the endowment (2 points) (see Supplementary Figure S6). The remaining clusters encapsulate less frequent behaviours (see Supplementary Figures S5, S7 and S8).
@@ III.3 Results for the model with Bayesian agents, data-based prior, and non-Bayesian agents
The results of the cluster analysis with N=100 return 4 clusters (86%, 6%, 2%, and 6%, respectively), shown in the Supplementary Figures S21 to S24. The behaviours are grouped into a main cluster that contains 86% of the simulations. These show agent behaviour that becomes more polarised as the rounds of the game progress towards highly polarised values (most agents donate all or nothing at the end of the game); see Fig. 9 as an example. Cluster 2 groups simulations with ultracooperative behaviour, i.e. those with an extremely high percentage of agents donating all the endowment (Fig. S22). The remaining clusters, which are less representative, show polarised simulations but to a lesser extent than the previous ones.
@@ III.4 Results for the model with Bayesian agents, non-informative prior, and non-Bayesian agents
The results of the cluster analysis with N=100 indicate the presence of six clusters, representing 82%, 5%, 9%, 1%, 1%, and 2%, respectively. Supplementary Figures S31 to S36 provide a visual representation of these clusters. The majority cluster, with 82% of the simulations, shows ultrapolarised behaviour, due to the presence of free riders and full cooperators (see Fig 11 and S31). With a non-informative prior, the dynamics of the agents are attracted by agents who do not change their opinion. The remaining clusters have simulations that are similar to cluster 1, with some particularities. They can be seen in Supplementary Figures S32 to S36.
Figure 11: Heatmap of the distribution of donations of a simulation of cluster type 1 behaviour of the model with Bayesian and non-Bayesian agents and non-informative prior (D) with N=100. The average donation is shown as dots joined with lines, in black.
The cluster analysis carried out with N=1000 delineates clusters with very similar instances, so we can consider that all behaviours have homogenised. Figure 12 and Supplementary Figure S37 visually illustrate the characteristics of this unique group.
Figure 12: Heatmap of the distribution of donations of a simulation of cluster type 1 behaviour of the model with Bayesian and non-Bayesian agents and non-informative prior (D) with N=1000. The average donation is shown as dots joined with lines, in black.
@@ IV Discussion and conclusions
In this study, we have introduced a model to elucidate human behaviour in the Public Goods Game when individuals possess more information than just the average contribution of the group, incorporating details on the distribution of donations from all participants in the previous round. The model assumption is that individuals in this scenario are not seeking to optimise their gains, but rather to converge and imitate the group’s behaviour. The model is implemented in four different versions (A to D), each with distinct parameterizations. The first model, model (A), is the simplest, assuming that individuals have no knowledge of how others behave in this game. Model (B) refines this assumption, incorporating some form of knowledge, drawing insights from real experimental data [10]. The last two models extend these initial models by introducing individuals who are not influenced by the group but exhibit independent behaviours such as free riders or full cooperators, as observed in experiments.
@@
Specifically, Model (C) with 100 agents exhibits behavioural patterns compatible with those experimentally observed in groups of 100 individuals (see Figures 1 and 1 vs 9 and Supplementary Figure S21). Here, we observe how agent behaviour and human behaviour in the initial rounds stem from a prior assumption and subsequently evolve, polarising towards situations with a high percentage of full cooperators and free riders.
These findings could be attributed not only to the effect of Bayesian agents but also to a small-scale effect. In fact, we note that for larger group sizes, as seen in the results with 1000 agents presented throughout the study, the agent behaviours become much more homogeneous across all models. This homogenisation of behaviours in large populations directly results from employing larger samples for Bayesian inference. Since individuals update their priors throughout the game’s evolution using observational data from the previous round, a larger population implies a larger data sample to estimate the true distribution of group decisions, and this larger sample expedites convergence to a steady state. This steady-state result to be the initial condition (plus the percentage of defectors and free riders when present).
In fact, on larger scales, such as with N = 1000 individuals, the model outcomes do not align with the experimental findings. This discrepancy arises because the experimental results from the sole replication involving 1000 participants resemble those observed with only 100 individuals. Consequently, drawing definitive conclusions and establishing causation proves challenging given the limited empirical evidence currently available. One potential explanation for this discrepancy may lie in individuals’ inability to effectively compute and estimate priors and posteriors in groups larger than those that human’s brains can handle, as suggested by Dunbar’s number [30], which is 150. The Dunbar’s number refers to a theoretical cognitive limit to the number of people with whom one can maintain stable social relationships. Beyond this limit, it becomes increasingly difficult for individuals to maintain meaningful relationships and track social dynamics effectively.
Our results suggest that humans, in Public Goods Game scenarios, might prefer imitation and convergence behaviours over profit optimisation, providing a plausible explanation for the surprising results, high levels of average cooperation and polarisation, in Public Goods Dilemmas of this kind (individuals with more information). Indeed, our results align with previous findings on group conformity. Herding behaviour has been shown to be influenced by several factors, namely payoff externalities, correlated effects, information externalities, and social preferences. Payoff externalities manifest when an individual’s actions impact the payoffs of others, leading to a collective choice and equilibrium [31, 32, 33]. Similarly, correlated effects arise when agents adopt similar behaviours due to shared external influences [34]. In our work, agents are exposed to information externalities, where previous works showed that individuals base their decisions on information inferred from observing the choices of their predecessors, leading them to disregard their private information. Consequently, their decisions become uninformative for subsequent players, initiating a cascade effect [35, 36, 37, 38, 39]. In the psychology literature, individuals are also assumed to have a preference for conformity [40, 41].
Finally, adhering to Occam’s Razor, our model, while limited, offer generic insights into how cooperation is achieved through the convergence to group behaviour. A potential extension of the model would be introducing noise or trembling hand in decisions, as found in [28], where ‘people are prone to make errors and random decisions’. Simulations with this added noise would study which level of noise is able to destroy cooperation and how cooperation (distribution and average) and group conformity are affected by different levels of trembling hand. This avenue presents an intriguing prospect for further exploration.